# BMG-Q: Localized Bipartite Match Graph Attention Q-Learning for Ride-Pooling Order Dispatch

### v1

@@ VII Conclusion
This paper proposes the Localized Bipartite Match Graph Attention Q-Learning (BMG-Q), a novel effective, scalable, and robust MARL algorithm framework tailored for large-scale ride-pooling order dispatch. By integrating localized bipartite match within the MDP of the ride-pooling system, we developed GATDDQN as a novel MARL backbone to accurately capture the dynamic interactions among agents in the large-scale ride-pooling order dispatch systems. Enhanced by gradient clipping and localized graph sampling, our GATDDQN improves scalability and robustness for very large-scale system, while the inclusion of a posterior score function in ILP captures the online exploration-exploitation trade-off and assists to reduce potential overestimation bias of agents. Through extensive experiments and validation, we show that BMG-Q demonstrates a superior performance in both training and operations of thousands of vehicle agents, outperforming benchmark RL frameworks by around 10% in accumulative rewards and showing a significant reduction in overestimation bias by over 50% while maintaining robustness and effectiveness amidst task variations and fleet size changes. Potential enhancements to our framework could be achieved by extending its application to multimodal/intermodal transportation systems [5, 4]. Additionally, refining the framework by integrating BMG-Q learning with KL-control methods [67] or conducting a more thorough theoretical analysis and proof of the underlying MDP may bring significant further advancements [58].

### v2

@@ VII Conclusion
This paper proposes the Localized Bipartite Match Graph Attention Q-Learning (BMG-Q), a novel effective, scalable, and robust MARL algorithm framework tailored for large-scale ride-pooling order dispatch. By integrating localized bipartite match within the MDP of the ride-pooling system, we developed GATDDQN as a novel MARL backbone to accurately capture the dynamic interactions among agents in the large-scale ride-pooling order dispatch systems. Enhanced by gradient clipping and localized graph sampling, our GATDDQN improves scalability and robustness for very large-scale system, while the inclusion of a posterior score function in ILP captures the online exploration-exploitation trade-off and assists to reduce potential overestimation bias of agents. Through extensive experiments and validation, we show that BMG-Q demonstrates a superior performance in both training and operations of thousands of vehicle agents, outperforming benchmark RL frameworks by around 10% in accumulative rewards and showing a significant reduction in overestimation bias by over 50% while maintaining robustness and effectiveness amidst task variations and fleet size changes. Future extensions of our BMG-Q framework could explore applications in multimodal transportation systems [14], power system operations [29], and other spatial-temporal decision games [44], leveraging its inherent strengths in handling diverse network topologies and multi-agent interdependence across these domains.